April 29, 2011

Data saved in /data/luci/20110430

Observers: Jill Bechtold, Olga Kuhn, Ian McGreer, Wiphu R., Linhua Jiang

TO: Steven

Afternoon: Very windy

6pm. Too windy to open. Dave Ashby is doing balance test. Dark frame looks ok on Luci

7pm. Still too windy to open.

10:35pm. Opening.

10:50pm. Collimating near SDSS1038+48. Gusty winds

10:55pm sdss1038.K.txt Ks. seeing 1.3" on guider. Very clear out.

start file luci.20110430.0003

end file luci.20110430.0012

11:30pm Started sdss1038.slit1.acq.txt

11:45pm Started sdss1038.spec.txt 1.3" seeing

200 H+K 1.88 microns Order Sep. LS600 slit

300s x 15 exposures

First file: luci.20110430.0018

Halpha! Cool!

12:20am seeing 1.57"

final file: luci.20110430.0032

Off to telluric. No star found -- pointing off. Steven is fixing pointing. Ambient temperature is dropping fast.

acquired tell.

1:20am. Started telluric script. file 36-37

Off to Linhua's 2nd MOS. 1324+27

2am. Started science exposures. Seeing 1" - 1.2"

First file 30 min. file 41.

Taking 2nd exposure.

1 degree gradient across the primary all night

3am off to sdss1420.K.txt

beachballs!

waiting for collimation

Seeing still blechy

3:10am SDSS1420+39 Ks files 43 - 52

3:35am SDSS1420_39 H files 53 - 62

4am. Doing CasA light echo in K files 63-77

Sky flats.

78-87 H-band, low

88-92 H-low

93-97 H High

98-102 Ks High

103-112 Ks Hi

113-122 Ks High

124-125 test

126-130 Ks, about 8000 counts

135-144 Ks about 10,000 counts

Closed dome, cleans and darks
Topic revision: r7 - 30 Apr 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback